OpenStudy (anonymous):

What is the third–degree polynomial function such that f(0) = –24 and whose zeros are 1, 2, and 3? A. f(x) = 4x^3 – 24x^2 + 44x – 24 B. f(x) = 2x^3 – 24x^2 + 22x – 24 C. f(x) = 4x^3 + 24x^2 – 44x – 24 D. f(x) = 2x^3 + 24x^2 – 22x – 24

OpenStudy (anonymous):

We write like this. y=(x-3)/3 and we need to get x rather than y. So y=x/3 -1 => x/3=y+1 => x=3y+3 So f(y)=3y+3. Or f-1(x)=3x+3. Hope it helped. Cheers!

OpenStudy (anonymous):

1. C 2. A 3. B

OpenStudy (anonymous):

1. [C] x^3 - 3x^2 - 5x + 15 2. [A] x^3 - 4x^2 + 4x - 16 3. [B] f(x) = 2x^3 - 12x^2 + 22x - 12 Good luck!

OpenStudy (anonymous):

doesn't make sense -.-

OpenStudy (anonymous):

LoL I know